OpenStudy (anonymous):

Find the next number in the number pattern 192, 96, 48, 24, ...

OpenStudy (anonymous):

12 i think

OpenStudy (anonymous):

i got 12

OpenStudy (anonymous):

i got 12 also 12+12=24+24=48+48=96+96=192

OpenStudy (anonymous):

yea haha i was wandering how to make sense of that answer but you get it xD.